# MGAug: Multimodal Geometric Augmentation in Latent Spaces of Image Deformations

### 20 Dec 2023 v1

@@ Introduction
Evaluation of segmentation on real 3D brain MRIs. We validate the performance of all methods on real 3D brain images with a decent size of training data. In this experiment, we employ TransUNet as our network backbone, given its proven superiority in brain segmentation performance (Chen et al. 2021). To ensure a fair comparison of the final segmentation accuracy, we augment the training data three times for all methods, resulting in a total number of approximately $2000$ augmented images. We then compare the segmentation dice scores of all augmentation methods over different brain structures.
Parameter setting. We set parameter $\alpha=3.0$ for the operator $L$ and the number of time steps for Euler integration in EPDiff (Eq. (3)) as $10$. We set the noise variance $\sigma$ as $0.02$, the batch size as $16$, and use the cosine annealing learning rate scheduler that starts from a learning rate $\eta=1e^{-3}$ for network training. We run all the models on training and validation images with Adam optimizer until obtaining the best validation performance. The training and prediction procedure of all learning-based methods are performed on a 40GB NVIDIA A100 Tensor Core GPU. For both 2D and 3D datasets, we split the images by using $70\%$ as training, $15\%$ as validation, and $15\%$ as testing images.
Results
The left panel of Fig. 3 reports our model MGAug’s classification accuracy (using ViT as the network backbone) on 2D silhouette shape dataset with different number of modes, $C$, along with diverse levels of augmented images quantities (i.e., 5×,7×,5$\times$,7$\times$,5 × , 7 × , and 10×10$\times$10 × the original training data size). It suggests that the utilization of a multimodal latent space $(C>1)$ consistently yields superior performance compared to an unimodal distribution $(C=1)$ by more than $2$-$7\%$. Our model achieves an optimal performance when $C=7$, which represents the number of classes in the training data. The right panel of Fig. 3 shows a comprehensive evaluation of classification accuracy across all augmentation methods with an increased number of augmented images in training data. Our method significantly outperforms other baselines by more than $7\%$ of accuracy with different numbers of augmented images. Another interesting observation is that our MGAug achieves comparable classification accuracy when using roughly 3×3$\times$3 × augmented images for training vs. other baselines using 10×10$\times$10 × of ground truth images. This indicates the potency of our method in efficiently leveraging augmented data to achieve competitive results.
Table 1: Accuracy comparison on 2D shapes over all models with various network backbones.
@@ Results
Tab. 1 reports the multi-class classification accuracy on a relatively smaller 2D silhouette shape data across all augmentation methods with various classification network backbones. For a fair comparison, we augment the data 10×10$\times$10 × of the original training images. MGAug achieves the highest performance, outperforming all baselines by $2$-$9\%$ of accuracy on six different network backbones.
@@
The left panel of Fig. 6 showcases the classification results on the 2D MNIST dataset for all methods, using different percentages of ground truth data during the training phase. Our MGAug method consistently achieves the highest accuracy across all augmentation techniques. Remarkably, MGAug reaches an accuracy of $98.02\%$ (which is the baseline accuracy without any augmentation) using only $20\%$ of the ground truth data, whereas other methods require at least $50\%$ of ground truth data to achieve similar results. Similarly, the right panel of Fig. 6 presents the dice evaluation for 3D brain MRI segmentations, as the percentage of ground truth data used in training is progressively increased. In a manner mirroring the classification results, our MGAug consistently outperforms the baseline algorithms by yielding the highest dice scores. These results indicate that our augmentation method has great potential to better alleviate the heavy need for labor and time-consuming ground-truths.
Figure 6: Left to right: classification evaluation on 2D MNIST dataset; segmentation evaluation on 3D brain MRIs with various percentages of ground truth used in the training inference. The horizontal dashed line (- -) serves as a reference, representing the performance achieved when utilizing all available training images without any augmentation.
@@
It is to our attention that diffusion models have recently become a powerful generative model (Ho, Jain, and Abbeel 2020). In this context, we initiate a preliminary comparison between our MGAug model and data augmentation through a diffusion model operating in the image space (by manipulating image intensities and textures) (Trabucco et al. 2023). This comparison was carried out with a specific focus on classification tasks using the 2D MNIST dataset.
@@
MGAug vs. intensity-based augmentations. While the primary emphasis of this paper does not fall into the realm of intensity-based augmentations, we compare MGAug with a few notable DA schemes that manipulate image intensities or textures. These methods include AutoAugment (Cubuk et al. 2019), RandAugment (Cubuk et al. 2020), AugMix (Hendrycks et al. 2019), and TrivialAugment (Müller and Hutter 2021).
Tab. 2 reports the accuracy comparison of a variety of DA schemes for image classification networks based on the ViT backbone. The testing results on both MNIST digits and 2D shape data show that our model consistently outperforms the baselines.
Table 2: Accuracy comparison on both 2D digits and 2D shape datasets over a mix of augmentation methods (including intensity-based vs. transformation-based DA) on ViT.
@@
Conclusion & Discussion
In this paper, we present a novel data augmentation model, MGAug, which learns augmenting transformations in a multimodal latent space of geometric deformations reparameterized by low-dimensional tangent spaces of diffeomorphisms. In contrast to existing augmentation models that adopt a unimodal distribution to augment data, our augmentation method fully characterizes the multimodal distribution of image transformations via introducing a mixture of multivariate Gaussians as the prior. Our method significantly improves the performance of multi-class classification (on 2D dataset) and image segmentation (on 3D brain MRIs) and eliminates the artifacts of geometric transformations by preserving topological structures of augmented images. While MGAug focuses on LDDMM, the theoretical tool developed in our model is widely applicable to various registration frameworks, for example, stationary velocity fields (Modat et al. 2012).
@@
Acknowledgments.
This work was supported by NSF CAREER Grant 2239977.

### 25 Jan 2024 v2

@@ IV-B Experiments and Implementation Details
TABLE I: Accuracy comparison on 2D shapes over all models with various network backbones.
@@
Evaluation of segmentation on real 3D brain MRIs. We validate the performance of all methods on real 3D brain images with a decent size of training data. In this experiment, we employ TransUNet as our network backbone, given its proven superiority in brain segmentation performance [8]. To ensure a fair comparison of the final segmentation accuracy, we augment the training data three times for all methods, resulting in a total number of approximately $2000$ augmented images. We then compare the segmentation dice scores of all augmentation methods over different brain structures.
Parameter setting. We set parameter $\alpha=3.0$ for the operator $\mathcal{L}$ and the number of time steps for Euler integration in EPDiff (Eq. (3)) as $10$. We set the noise variance $\sigma$ as $0.02$, the batch size as $16$, and use the cosine annealing learning rate scheduler that starts from a learning rate $\eta=1e^{-3}$ for network training. We run all the models on training and validation images with Adam optimizer until obtaining the best validation performance. The training and prediction procedure of all learning-based methods are performed on a 40GB NVIDIA A100 Tensor Core GPUs. For both 2D and 3D datasets, we split the images by using $70\%$ as training, $15\%$ as validation, and $15\%$ as testing images.
IV-C Results
Tab. I reports the multi-class classification accuracy on a relatively smaller 2D silhouette shape data across all augmentation methods. For a fair comparison, we augment the data 10×10$\times$10 × of the original training images. MGAug achieves the highest performance, outperforming all baselines by $2$-$9\%$ of accuracy on six different network backbones.
The left panel of Fig. 3 reports our model MGAug’s classification accuracy (using ViT as the network backbone) on 2D silhouette shape dataset with different number of modes, $C$, along with diverse levels of augmented images quantities (i.e., 5×,7×,5$\times$,7$\times$,5 × , 7 × , and 10×10$\times$10 × the original training data size). It suggests that the utilization of a multimodal latent space $(C>1)$ consistently yields superior performance compared to an unimodal distribution $(C=1)$ by more than $2$-$7\%$. Our model achieves an optimal performance when $C=7$, which represents the number of classes in the training data.
The right panel of Fig. 3 shows a comprehensive evaluation of classification accuracy across all augmentation methods with an increased number of augmented images in training data. Our method significantly outperforms other baselines by more than $7\%$ of accuracy with different numbers of augmented images. Another interesting observation is that our MGAug achieves comparable classification accuracy when using roughly 3×3$\times$3 × augmented images for training vs. other baselines using 10×10$\times$10 × of ground truth images. This indicates the potency of our method in efficiently leveraging augmented data to achieve competitive results.
@@ IV-C Results
The left panel of Fig. 6 showcases the classification results on the 2D MNIST dataset for all methods, using different percentages of ground truth data during the training phase. Our MGAug method consistently achieves the highest accuracy across all augmentation techniques. Remarkably, MGAug reaches an accuracy of $98.02\%$ (which is the baseline accuracy without any augmentation) using only $20\%$ of the ground truth data, whereas other methods require at least $50\%$ of ground truth data to achieve similar results. Similarly, the right panel of Fig. 6 presents the dice evaluation for 3D brain MRI segmentations, as the percentage of ground truth data used in training is progressively increased. In a manner mirroring the classification results, our MGAug consistently outperforms the baseline algorithms by yielding the highest dice scores. These results indicate that our augmentation method has great potential to better alleviate the heavy need for labor and time-consuming ground-truth labels.
Figure 6: Left to right: classification evaluation on 2D MNIST dataset; segmentation evaluation on 3D brain MRIs with various percentages of ground truth used in the training inference. The horizontal dashed line (- -) serves as a reference, representing the classification accuracy or dice score achieved when utilizing all available training images without any augmentation.
@@
It is to our attention that diffusion models have recently become a powerful generative model [24]. In this context, we initiated a preliminary comparison between our MGAug model and data augmentation through a diffusion model operating in the image space (by manipulating image intensities and textures) [56]. This comparison was carried out with a specific focus on classification tasks using the 2D MNIST dataset.
@@
MGAug vs. intensity-based augmentations. While the primary emphasis of this paper does not fall into the real of intensity-based augmentations, we compare MGAug with a few notable DA schemes manipulationg image intensities or textures. These methods include AutoAugment, an automated scheme that searches for improved augmentation policies [11]; RandAugment, an augmentation algorithm which reduces the computational expense by a simplified search space [12]; AugMix, a data processing method to improve the robustness and uncertainty estimates [22]; and TrivialAugment, a parameter-free data augmentation model [44].
Tab. II reports the accuracy comparison of a variety of DA schemes for image classification networks based on the ViT backbone. The testing results on both MNIST digits and 2D shape data show that our model outperforms the baselines consistently.
TABLE II: Accuracy comparison on both 2D digits and 2D shape datasets over a mix of augmentation methods (including intensity-based vs. transformation-based DA) on ViT backbone.
@@
V Conclusion
In this paper, we present a novel data augmentation model, MGAug, which learns augmenting transformations in a multimodal latent space of geometric deformations reparameterized by low-dimensional tangent spaces of diffeomorphisms. In contrast to existing augmentation models that adopt a unimodal distribution to augment data, our augmentation method fully characterizes the multimodal distribution of image transformations via introducing a mixture of multivariate Gaussians as the prior. Our method significantly improves the performance of multi-class classification (on 2D datasets) and image segmentation (on 3D brain MRIs) and eliminates the artifacts of geometric transformations by preserving the topological structures of augmented images. While MGAug focuses on LDDMM, the theoretical tool developed in our model is widely applicable to various registration frameworks, for example, stationary velocity fields [41, 19].